OpenStudy (anonymous):

tell me the way how to solve it througly : 1.2*10<2*2.8*10<4/4*1.4

OpenStudy (zepp):

That's arithmetic D:

OpenStudy (anonymous):

i'am weak in this area.

OpenStudy (anonymous):

and the answer is : \[6\times10<7 m\]

OpenStudy (anonymous):

Can you rewrite it properly.. what is dot(.) means there in between 2*2.8*10 ??

OpenStudy (anonymous):

Use equation editor or draw tool to write that..

OpenStudy (anonymous):

this one.what i was asking. i know it answer. but hav'nt understood how this calculation takes place plesae helppp

OpenStudy (anonymous):

And what you have written there is totally different..

OpenStudy (anonymous):

yes. i hav'nt used the editor well. now please tell me

OpenStudy (anonymous):

Cancel the terms 1 by 1..

OpenStudy (anonymous):

can u please use the equation editor to tell me please.?

OpenStudy (anonymous):

Yes I will..

OpenStudy (anonymous):

\[\large \frac{1.2 \times 10^{-2} \times 0.28 \times 10^{-3}}{4 \times 1.4}\]

OpenStudy (anonymous):

now what to do?

OpenStudy (anonymous):

See 1.2 can be written as: \(\frac{1}{2}\) 0.28 can be as: \(\frac{28}{100}\) 1.4 can be as : \(\frac{14}{10}\) \(10^{-2}\) can be : \( \frac{1}{100}\) \(10^{-3}\) can be : \( \frac{1}{1000}\)

OpenStudy (anonymous):

Sorry mistake there: 1.2 can be written as : \(\frac{12}{10}\)

OpenStudy (anonymous):

is the denominator as zero is given according to the inverse given?

OpenStudy (anonymous):

No...

OpenStudy (anonymous):

In case of multiplication when you cancel any terms then result is 1 not 0..

OpenStudy (anonymous):

then how you have putten the zero's in dnominator?

OpenStudy (anonymous):

See in case of decimals: \[x.y = \frac{xy}{10}\] After dot there is one term so we put 10 in denominator.. Getting??

OpenStudy (anonymous):

& if there is x.2y then is that be =x2y/100 ?

OpenStudy (anonymous):

similary: for: \[x.yz = \frac{xyz}{100}\] There is yz after dot so we will put 100..

OpenStudy (anonymous):

Yes absolutely correct..

OpenStudy (anonymous):

okay.thanks. now how shall we contineu?

OpenStudy (anonymous):

Yes..

OpenStudy (anonymous):

And remember: \[\large x^{-a} = \frac{1}{x^a}\] this is the case with \(10^{-2}\) and \(10^{-3}\) Getting??

OpenStudy (anonymous):

So what we get I show you after putting...

OpenStudy (anonymous):

1/10squr. and 1/10cube?

OpenStudy (anonymous):

\[\large \implies \frac{\frac{12}{10} \times \frac{1}{100} \times \frac{28}{100} \times \frac{1}{1000}}{4 \times \frac{14}{10}}\] Getting??

OpenStudy (anonymous):

yes..now please contineu

OpenStudy (anonymous):

See now you can bring 10 0r 100 or 1000 which are in numerator to the denominator and which are in denominator can be brought back to numerator.. Like this: \[\large \implies \frac{12 \times 28 \times 10}{ 10 \times 100 \times 100 \times 1000 \times 4 \times 14}\]

OpenStudy (anonymous):

which terms from denumerator are placed in numeraor?

OpenStudy (anonymous):

10..

OpenStudy (anonymous):

there is 14/10 so bring the 10 above see I have multiplied in the last in the numerator.. Getting??

OpenStudy (anonymous):

we will leave 4 and 14 still on denom.. tr?

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

Cancel what ever is being cancelled: \[\large \implies \frac{6 \times \cancel{10}}{10000000 \cancel0} \implies \frac{6}{10000000}\]

OpenStudy (anonymous):

See the terms that are denominator of denominator can be brought to numerator..

OpenStudy (anonymous):

in denominator 10 is again denominator so bring it up..

OpenStudy (anonymous):

Or you can remember it as: \[\large \frac{\frac{a}{b}}{\frac{c}{d}} = \frac{a \times d}{b \times c}\]

OpenStudy (anonymous):

how u brought 6/10000000

OpenStudy (anonymous):

Now back: \[\implies \frac{6}{10000000} \implies \frac{6}{10^7}\]

OpenStudy (anonymous):

you got its previous step of not?? \[\implies \frac{6 \times 10}{100000000}\] Is it clear to you??

OpenStudy (anonymous):

yes. but how it came?

OpenStudy (anonymous):